OpenStudy (ebonyserrano):

Write the equation of the line that passes through (1, 3) and has a slope of 2 in point-slope form. y – 1 = 2(x – 3) y – 3 = 2(x – 1) x – 1 = 2(y – 3) x – 3 = 2(y – 1)

OpenStudy (ebonyserrano):

@AloneS

jimthompson5910 (jim_thompson5910):

Similar example: Write the equation of the line that passes through (5,7) and has a slope of 9 Given slope = \(\Large {\color{red}{m}}={\color{red}{9}}\) Given point = \(\Large ({\color{blue}{x_1}},{\color{green}{y_1}})=({\color{blue}{5}},{\color{green}{7})}={\color{green}{7}}\) ---------------------------- You'll plug this given information into the point slope equation \[\Large y - y_1 = m(x - x_1)\] \[\Large y - {\color{green}{y_1}} = {\color{red}{m}}(x - {\color{blue}{x_1}})\] \[\Large y - {\color{green}{7}} = {\color{red}{9}}(x - {\color{blue}{5}})\]

alones (alones):

@Ebonyserrano Do youeven understand?

OpenStudy (ebonyserrano):

No I'm stupid

alones (alones):

Kk don't say that... Only noobs are stupid.

OpenStudy (ebonyserrano):

Noob is my middle name

alones (alones):

Oh sorry.

OpenStudy (mathmale):

The point-slope form of the equation of a str. line is y-k=m(x-h). The given point is represented by (h,k), and you are given the values of h and k, along with the value of the slope. Please plug h, k and m into the equation I've given you.

OpenStudy (mathmale):

jim's example is a good one. Study it until you understand what he is doing, and then apply what you have learned to solving the problem you were given.